The number of fish Jim catches is dependent upon the lures he uses. He also loses lots of these expensive lures. Here in South Dakota each fisherman can use two poles at one time so Jim can use two different lures. If L1 represents the number of fish caught per hour for the lure on the pole that holds and L2 represents the number of fish caught per hour for the lure on the pole that he puts in the "dummy" holder, then the formula below represents the relation between the two lures used and the total number of lures lost per hour using the two poles. (C is the rate of lures lost per hour using both poles.) If the combined loss rate and the rate for fish on lure 2 are known, write a function for the rate of fish on the first lure (i.e. solve the equation below for L1.)
1/C=1/L_1 +L_2/L_1

Answer by stanbon(48575)

solve the equation below for L1.)
1/C=1/L1 +L2/L1
-----------------------
Multiply thru by L1 to get:
L1/C = 1 + L2
------------------
Multiply thru by C to get:
L1 = C + C*L2
---------------------------
Cheers,
Stan H.